OpenStudy (anonymous):

solve for x : x-4+ 3/x=0 any ideas? i came to the solutions (1,3).

OpenStudy (campbell_st):

rewrite the equation x - 4 = -3/x multply each term by x x^2 - 4x = -3 add 3 to both sides x^2 - 4x + 3=0 factorise (x -3)(x+1) x = 3, -1

OpenStudy (anonymous):

@campbell_st I think it should be x - 4 = -3/x you missed the sign. So the solution is, x^2 -4x = -3 or, x^2 -4x +3 =0 or, x^2 -3x -x +3 =0 or, x(x-3) -1 (x-3) =0 so, x = 1,3 @brandon871861 you are right.

OpenStudy (campbell_st):

ummm thats what my solution is... please check

OpenStudy (anonymous):

thank you very much yall.. greatly appriciate it!!

OpenStudy (campbell_st):

the solution isn't (1,3) as this indicates an ordered pair... the solution is x = 1, 3

OpenStudy (anonymous):

i dont understand when you say x=1, 3

OpenStudy (anonymous):

is it the same things as saying... x=1 and x=3? i dont see how x can be both 1 and 3 if thats the case..

OpenStudy (campbell_st):

yes when you substitute x = 1 or x = 3 into the equation the solution is 0... the ordered pairs are (1, 0) and (3, 0) your solution (1, 3) is x = 1 and y = 3 the equation you solved can be drawn on the number plane... and the solutions x=1 and x = 3 is where it cuts the x axis

OpenStudy (campbell_st):

this is a simple graph